# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ II Tools and Definitions
This can be interpreted as the minimum amount of random bits a variable $X_{i}$ has, when conditioned on all previous random variables, indexed by $10,\ldots,i-$ 1.
Definition 3 (Min-entropy rate).
The min-entropy rate of a random variable $X\in\{0,1\}^{n}$ is
$\alpha=\frac{\mathrm{H}_{\infty}(X)}{n}.$ (3)
This can be interpreted as the minimum amount of randomness $X$ has per bit, on average.
Definition 4 (Statistical distance).
The statistical distance, $\Delta$, between two random variables, $X,Z\in\{0,1\}^{n}$ is defined as
This is a measure of how close, or indistinguishable, two random variables are to one another.
Definition 5 ($\epsilon$-perfect randomness).
A random variable $X$ on $\{0,1\}^{n}$ is said to be $\epsilon$-perfectly random, if,
$\Delta(X,\mathrm{U}_{n})\leq\epsilon,$ (5)
where $\mathrm{U}_{n}$ is the uniform variable on $\{0,1\}^{n}$, i.e. $\Pr(\mathrm{U}_{n}=u)=\frac{1}{2^{n}}$ for all $u\in\{0,1\}^{n}$.
This definition is equivalent to saying that the variable $X$ is distinguishable from a uniform distribution with distinguishing advantage at most $\epsilon$, i.e. the distinguisher in the game described above is successful with probability at most $\frac{1}{2}+\epsilon$. When $\epsilon=0$, the random variable is said to be perfectly random. This definition is universally-composable canetti2001universally , i.e. $X$ can be used safely in other applications.
III Statistical Testing
Statistical test suites are collections of algorithms that analyse the numerical properties of a set of random numbers to determine whether there is evidence to reject the possibility that they are uniformly distributed. If there is sufficient evidence to reject this possibility, a statistical test is said to be failed, which directly implies that the output can be distinguished from the ideal uniform distribution at some confidence level, as described in the previous section. The hypothesis that a random variable is uniformly distributed is known as the null hypothesis $\mathsf{H_{0}}$. For an RNG producing a random variable $X\in\{0,1\}^{n}$, the null hypothesis is $\mathsf{H_{0}}:\Delta(X,\mathrm{U}_{n})=0$. If the null hypothesis is rejected, then the alternative hypothesis $\mathsf{H_{1}}:\Delta(X,\mathrm{U}_{n})>0$ is accepted.
However, a random variable cannot be tested directly, only its realisation can – i.e. the bit string $x\in\{0,1\}^{n}$ produced by the random variable $X$. To assess whether to accept or reject the null hypothesis, a statistical test calculates a specific measure of $x$ (e.g., its mean), known as the test statistic $t$, and analyses how likely this test statistic is to be observed, assuming that the underlying random variable is uniform. Test statistics calculated from realisations of a uniform distribution are normally distributed, so one can calculate how likely observing certain ranges of the test statistic is by using concentration inequalities.
More precisely, this likelihood is captured by a probability known as the p-value, which can be defined as follows:
Definition 6 (p-value).
Given an observed test statistic $t$ obtained by calculating a measure from the realisation of a random variable $X=x\in\{0,1\}^{n}$ and $T$, the (normally distributed) variable associated with all the possible measure values, the p-value $\mathsf{p}\in[0,1]$ is defined as
$\mathsf{p}=\Pr(T\leq t|\Delta(X,\mathrm{U}_{n})=0)$ (6)
where $\mathrm{U}_{n}\in\{0,1\}^{n}$ is uniformly distributed.
A range of p-values is defined that provide a threshold at which the null hypothesis is rejected i.e. when the test is deemed to fail. If a test ensures there is, at most, a $1\%$ chance it incorrectly rejects that the RNG is producing uniform random numbers [^4] Known as the type 1 error – when a statistical test incorrectly rejects a true null hypothesis., then it would, for example, conclude failure if $\mathsf{p}\notin[0.01,1]$. This threshold for failure is on one tail only, so it only fails test statistics that are sufficiently biased away from the expected value in one direction. More generally, tests are two-tailed and conclude failure if the observed p-values are outside of a sufficiently large interval, for example, if $\mathsf{p}\notin[0.005,0.995]$.
The failure of numerous statistical tests is a strong indicator that an RNG is not producing (near-)perfect random numbers, as its output can be distinguished from the uniform distribution with high probability. For example, if all statistical test performed on the RNG are independent, the probability that the null hypothesis is accepted given that the alternative hypothesis is true (known as the type 2 error) is $\mathsf{p_{type2}}=\mathsf{p^{test1}_{type2}}\cdot\mathsf{p^{test2}_{type2}}\cdot\ldots\cdot\mathsf{p^{testn}_{type2}}$, where $n$ is the number of tests performed.
III.1 Existing test suites
III.1.1 NIST statistical test suite
The NIST statistical test suite (SP 800-22) rukhin2001statistical is the best known and widely used. This suite contains 15 tests, some of which have multiple sub-tests, and passing them is a requirement for RNG certification by numerous governing bodies such as NIST and BSI. During testing, a file of randomness is split into sub-strings, and each sub-string is tested individually. The user can define the number of sub-strings and the total bit string size to analyse.
The user guide suggests using 100 sub-strings of $10^{6}$ bits, which requires a minimum of $10^{8}$ bits or equivalently, 12.5 megabytes (MB), for testing. For each statistical test, a set of p-values is calculated, where each p-value corresponds to one of the sub-strings. The pass/fail analysis is then performed using this set of p-values, giving two test results. The first result is a statistical test on the observed p-values, assessing the null hypothesis that the set of p-values is uniformly distributed, at the $1\%$ significance level. The second result is to check that sufficiently many sub-strings pass the test at the $1\%$ confidence level (i.e. have p-values in the range $[0.005,0.995]$). In order for an RNG to be deemed as producing satisfactory random numbers, it must pass both results for each test. See mckay2016users for further details.
III.1.2 Diehard(er) statistical test suite
The Dieharder statistical test suite consists of the 18 Diehard tests and additional tests, including some from the NIST test suite. Similarly to the NIST suite, this is one of the core test suites used by RNG certification bodies. Failure is concluded when $\mathsf{p}\notin[0.0005,1-0.0005]$ and a test is deemed ‘weak’ if $\mathsf{p}\in[0.0005,0.005]\cup[1-0.0005,1-0.005]$. This high tolerance for poor test statistics means that a bad RNG may sometimes pass the Diehard(er) test suite, but failing Diehard(er) is a strong indicator of non-uniformity.
The Diehard(er) tests require a significant amount of random numbers to avoid re-use of the input random numbers, which gives erroneous results. For this reason, we suggest test sizes of at least $1$ gigabyte (GB) of random numbers. If testing smaller file sizes, one can modify the default parameters to avoid these issues. Our testing environment uses the default parameters for each test.
III.1.3 TestU01 statistical test suite
TestU01 is a software library written in C that conducts RNG statistical testing with pre-compiled statistical test batteries. These batteries vary widely in the amount of tests and the amount of randomness they require. For full details of which tests these batteries include, see l2007testu01 .
Test p-values are shown if $\mathsf{p}\notin[0.001,0.999]$, so we will use this as our failure criteria. During our testing, we use the Alphabit, Rabbit and SmallCrush batteries, which are all contained in TestU01. In order to run these tests, files should contain at least $2^{25}$ random bytes ($\approx 35$MB). We omit Crush and BigCrush from our work due to their excessive runtime and their large random number requirement, however they can be executed with our statistical testing environment.
III.1.4 ENT statistical test suite
The ENT test suite is a small but efficient set of 6 statistical tests. This test suite has been used to show bias in a commercial quantum RNG by showing consistent failure in the so-called $\chi^{2}$ test hurley2017certifiably (we replicate these results of an RNG which we acquired independently, see Table 3).
The ENT tests output test statistics directly, without giving a pass/failure threshold, so we assess failure based on the table found in Table 3 of ortiz2018heartbeats . Although there is no specific guidance on required input lengths, we found that the tests give suspicious results when input sizes are below $0.5$GB.
III.1.5 PractRand statistical test suite
PractRand is a C++  library of statistical tests for RNGs. It was designed with practicality in mind – to be efficient, user-friendly and detect significant bias in RNGs. According to its documentation, it boasts quicker runtime than most test suites (which we confirm, see Table 2), unique interfacing, no (in principle) maximum input length limit and some original tests. It performs tests based on size of the input file, testing on subsets of size $2^{24+x}$ bytes for $x\in\mathbb{N}$, and performing more tests as $x$ increases. In our testing, we limit the maximum test size to $2^{32}$ bytes ($\approx 4.3$GB). For more information, see doty2018practrand , where they have full details and additional analysis, including comparisons between PractRand and other test suites.
There are numerous result ranges that p-values may enter when testing using the PractRand suite, these are “unusual’, “mildly suspicious”, “suspicious”, “suspect” and “fail”. Failure is concluded when $\mathsf{p}\notin[10^{-11},1-10^{-11}]$.
III.2 Our statistical testing environment
The interfacing code for our $\mathsf{STE}$ can be downloaded at https://github.com/CQCL/random_test. We provide a Light, Recommended and All setting for statistical testing which is detailed below and can be executed using run_light, run_recommended and run_all commands respectively. The NIST statistical test suite is not performed with these commands, since it requires user prompts, but it can be executed individually in the environment. We believe that the recommended setting offers a nice trade-off between computational (and thus, environmental) cost and rigorousness, yet goes beyond standard numerical testing required by certification bodies. Using the $\mathsf{STE}$ (or by downloading, parameterising and executing the relevant statistical test suites independently), all results in this work can be replicated.
III.3 Suggested settings
We now suggest settings for statistical testing, based on knowledge acquired during this research. The runtimes shown are averaged over 10 executions and result from testing a 10Gbit file, except for NIST, where a 100Mbit file is tested (to align with the user guidance). Statistical testing was run on a Dell Precision 7540 personal laptop with 16GB of RAM and a 2.3GHz Intel I9-9880H processor, using the Ubuntu 20.04 operating system.
Test Mode
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
Total
Runtime
Total
Tests
Light Y Y Y 4m 44s 941
Recommended Y Y Y Y Y 114m 31s 999
All Y Y Y Y Y Y Y 127m 41s 1015
Table 1: This table details our settings for light, recommended and all statistical testing using the code provided. A ‘Y’ in a specific column indicates that the associated test suite of that column is included in the setting.
@@ III.3 Suggested settings
Table 2: This table gives the average runtime of all statistical test suites contained in our statistical testing environment. The average is taken running each test suite 10 times on independent inputs. For the NIST test suite, the runtime relates to testing a 100Mbit file. For all other test suites, the runtime is for a 10Gbit file.
III.3.1 Light
@@
Our suggested light test mode (executed with terminal command run_light) of ENT, SmallCrush and PractRand takes under 5 minutes to run and consists of approximately 941 tests. In our numerical testing, we show that this is a sufficient set of tests to detect failure to generate uniform randomness in the RNGs that fail (see Section IV and Section V).
III.3.2 Recommended
@@
Our recommended setting (executed with terminal command run_recommended) includes most light mode suites, with the exception of exchanging TestU01 SmallCrush for Rabbit, but also adding the NIST and Diehard tests. The inclusion of these additional tests cause the runtime to become around 2 hours and increase the total number of tests to 999. The only reduction from the full statistical test mode (executed with terminal command run_all) is the exclusion of TestU01 SmallCrush and Alphabit, which are omitted since most of the individual tests are highly correlated to those in Rabbit, for example, the same test with slightly different parameters. The recommended test suite includes all statistical tests required by RNG governing bodies (e.g. NIST and Dieharder), whilst being significantly more powerful than running those alone. In later sections, we show the importance of going beyond these individual test suites, in the sense that we show that an RNG that passes the NIST and Dieharder tests shows significant statistical bias once analysed with our combined STE (see Section IV).
III.4 Shortcomings of statistical testing
@@
Fundamentally, statistical tests have limited ability to validate if good random numbers are being produced by an RNG. They should rather be understood as a useful tool to detect failure to generate uniform random numbers, since passing statistical tests gives no guarantee of (near-)perfect randomness. This is especially important in the case of cryptographic RNGs. For example, in shrimpton2015provable , a thorough analysis of Intel’s RDSEED hardware RNG is performed and one of their conclusions is that “RDSEED delivers truly random bits but with a security margin that becomes worrisome if an adversary can see a large number of outputs from either interface. If he controls an unprivileged process on the same physical machine, this could happen very quickly” [^5] In this case, the adversary also requires control of an “unprivileged” process, which is a form of side-information that may be hard to obtain in practice. (on page 4). As we shall see next, our statistical testing results do not detect that RDSEED’s output can be distinguished from uniform.
@@
At the implementation level, the available software for numerous statistical test suites have been shown to have issues. For the NIST test suite, the list of implementation issues is extensive, so we summarize a few problems that the reader may find interesting. Research has found significant dependencies between the tests burciu2019systematic and implementation issues with certain tests; for example hamano2007correction found that the settings of both the Discrete Fourier Transform test and Lempel-Ziv test were wrong and kowalska2022revision found an error within the probability calculations for the Overlapping Template Matching test. Moreover, problems with how results are analysed were discovered, for example, marton2015interpretation found that although the NIST documentation provides guidance that the analysed RNG is random if all tests are passed, even though truly uniform data has a high probability (80%) of failing at least one NIST statistical (sub-)test. Some work even suggested that the tests are “harmful” saarinennist , namely that “The weakest pseudo-random number generators will easily pass these tests, promoting false confidence in insecure systems.”. During this work, we found an additional issue with the NIST test suite that we could not find reported elsewhere: the results showed all tests failed whenever the CPU was being used for other computations simultaneously. The NIST Random Bit Generation Team have been made aware of this. Other test suites have also had their own reported problems, including Dieharder. In sys2022bad they find that over 50% of the Dieharder tests generated biased null hypothesis distributions (which are expected to be uniform).
IV Statistical Testing of Different RNGs
@@
In this section, we use our $\mathsf{STE}$ to analyse the statistical properties of random numbers produced by some commonly used RNGs. At this stage, we do not apply any post-processing to the RNGs output, however, some of the RNGs that we consider already have post-processing included, in the form of so-called conditioners or deterministic randomness extractors. Therefore, in those cases our statistical analysis applies to the joint system comprising both the source of randomness and the existing post-processing in the device. We then also use and discuss a NIST min-entropy estimation tool, which provides a min-entropy estimate for use in the second part of our work in which we add and analyse further post-processing.
The RNGs that we analyse are:
  • 32-bit LFSR: a software pseudo-RNG.
@@
  • Intel RDSEED jun1999intel : a hardware RNG based on thermal noise from a ring oscillator, i.e. a chaotic process.
@@
  • IDQuantique (IDQ) Quantis quantique2004quantis : a hardware RNG based on the quantum effect of detecting photons at the output of a semi-transparent mirror.
@@
Further details and description of the RNGs can be found in Section A.3. The LFSR is still widely used in numerical simulations, although it is known to have flaws stkepien2013statistical . In our work, we study it mainly as a way to benchmark against, serving as the obvious bad choice in the sense that its output exhibits patterns – but we will see that the effect of post-processing its output is statistically non-trivial. Both Intel’s RDSEED and IDQ’s Quantis are sold as RNGs for cryptographic use. These RNGs are used in numerous applications and are a sample of the different types of RNGs available today. The statistical analysis is performed using the run_all function in our statistical test environment on $10\times 10$Gbit files from each RNG and, similarly, using the NIST test suite performed on $10\times 100$Mbit files split into 100 sub-strings each of $1$Mbit. The NIST min-entropy estimators mckay2016users are used in the so-called non-IID setting on $10\times 1$Mbit files. This analysis far exceeds that required by certification bodies, so may be a result of independent interest. All testing is done using the default parameters, unless otherwise stated.
IV.1 Results
@@
An RNG producing near-perfect randomness should pass almost all statistical tests. More concretely, we mean that the ideal RNG would fail less than 7.5 of the 4600 individual statistical tests [^6] This number is the expected amount of type 2 error, i.e. the expected maximum number of failed tests, given that the underlying distribution is indistinguishable from uniform. Note that we are implicitly assuming each statistical test is independent., on average. The results we obtained for the three RNGs, are summarised in the following Table 3 and displayed visually in Figure 4 (level 0).
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
32-bit LFSR 10 40 (3) 5 51 73 131 855 (167)
RDSEED 0 0 (4) 0 1 0 1 0 (7)
IDQ Quantis 0 0 (3) 5 0 17 25 3 (15)
@@
Table 3: This table gives the average sum of tests failed for $5\times 10$Gbit samples from each RNG (after testing 10 samples). The results are presented in this way to allow for direct comparison to later results, where only $5\times 10$Gbit samples are tested. Due to the 32-bit LFSR failing PractRand quickly, only 635 tests were conducted (instead of the full 4600) so we rescale these results. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. The full results can be found in Appendix B.
@@
In the statistical tests, the RDSEED RNG performs best, failing the least tests and, without surprise, the 32-bit LFSR performs worst. The poor performance of the LFSR is likely due to its periodicity, since bits repeat every $12^{32}-$ 1 ($4.3$Gbits) and this is less than the size of the files tested. The IDQ Quantis device performs well in the NIST and Diehard tests however fails an ENT test and several of the TestU01 suites tests. These observations reproduce (and add confidence to) the results of previous work hurley2020quantum . These results, especially for IDQ’s device, exhibit the need to go beyond the requirements of certification bodies for statistical testing, with additional tests providing a noticeable advantage in detecting failures.
@@
The NIST min-entropy estimators mckay2016users are a collection of algorithms that give a standardised way of estimating the min-entropy (as defined in Definition 1) of an RNG’s output. These estimators are both useful to evaluate the entropy generation of the studied RNG and to calculate a min-entropy bound that we use later to choose the randomness extractors parameters.
@@
Table 4: This table shows the average NIST min-entropy estimator, the sample standard deviation and a lower bound for min-entropy/bit for each RNG. These results are the average of 10 tests on different 1’000’000 bit samples, each generated with significant time gaps between the generation of different output test samples. Full results tables can be found in Section B.2.
In Table 4, the NIST min-entropy estimator per byte is the average observed per-byte min-entropy calculated by the NIST min-entropy estimator tool and $\overline{\mathsf{est}}$ is the per-bit average. The sample standard deviation, $\sigma$, describes how much the different test results fluctuated, which we calculate using the expression Equation 18 and $\alpha$ is a lower bound (with probability at-least $1-2^{-32}$) on the per-bit min entropy for any test sample. Details of this derivation can be found in Section B.2.1.
The estimated min-entropy per bit for the IDQ Quantis device was the largest (0.895). In terms of sample standard deviation, the 32-bit LFSR was by far the highest, which indicates the observed min-entropy estimators for different samples fluctuated the most. NIST recommends that a good RNG should have a of min-entropy per bit of at least $1-2^{-32}$ buller2022discussion , which is much larger than the values we observe from using their min-entropy estimator tool. That being said, results suggest that some estimator tests provide significant underestimates zhu2017analysis , which could explain the large disparity between the estimated results and the recommendation of NIST – but underestimates are not a problem in our case.
V A Variety of Post-processing Methods
V.1 Overview
Randomness extractors are mathematical algorithms that distil weakly random bit strings [^7] More precisely, a necessary (but not sufficient) condition for randomness extraction to be successful is that the source has some min-entropy, see Definition 1., in the sense that they are not uniformly distributed, into a near-perfect random bit string. In this section, we present, implement and test a variety of randomness extraction processes. The main question that we want to answer is whether these methods have an observed impact on the statistical properties of the RNGs output. The recipe that we follow is the following:

### v2

@@ 3.2.1 Suggested Settings
Table 2: This table gives the average runtimes of all statistical test suites contained in our statistical testing environment. The average is taken when running each test suite 10 times on independent inputs. For the NIST test suite, the runtime relates to testing a 100 Mbit file. For all other test suites, the runtime is for a 10 Gbit file.
@@
Light
@@
Our suggested light test mode, executed with the command run_light, includes ENT, SmallCrush, and PractRand, running in under 5 min and covering approximately 941 tests. Our numerical analysis shows that this set of tests is sufficient to detect failures in generating uniform randomness for RNGs that fail (see Sections 4 and 5).
Recommended
@@
Our recommended setting, executed with the command run_recommended, includes most of the light mode test suites, replacing TestU01 SmallCrush with Rabbit and adding the NIST and Diehard tests. These additional tests increase the runtime to approximately 2 h and bring the total number of tests to 999. The full statistical test mode (run_all) includes all suites, except for SmallCrush and Alphabit, which are omitted due to their significant correlations with the Rabbit tests, where only the parameters differ slightly. The recommended suite covers all tests required by RNG certification bodies (e.g., NIST and Dieharder) while offering a more comprehensive analysis than when running these tests alone. The later sections demonstrate the value of extending beyond individual test suites, as an RNG that passes the NIST and Dieharder tests can still show significant statistical bias when analysed with our combined STE (see Section 4).
3.3 Shortcomings of Statistical Testing
@@
Fundamentally, statistical tests have a limited ability to validate that good random numbers are being produced by an RNG. They should rather be understood as a useful tool to detect a failure to generate uniform random numbers, since passing statistical tests gives no guarantee of (near-)perfect randomness. This is especially important in the case of cryptographic RNGs. For example, in Shrimpton and Terashima (2015), a thorough analysis of Intel’s RDSEED hardware RNG is performed, and one of their conclusions is that “RDSEED delivers truly random bits but with a security margin that becomes worrisome if an adversary can see a large number of outputs from either interface. If he controls an unprivileged process on the same physical machine, this could happen very quickly” (in this case, the adversary also requires control of an “unprivileged” process, which is a form of side information that may be difficult to obtain in practice) (on page 4). As we shall see next, our statistical testing results do not detect that RDSEED’s output can be distinguished from uniform.
@@
At the implementation level, the available software for numerous statistical test suites has been shown to have issues. For the NIST test suite, the list of implementation issues is extensive, so we summarise a few problems that the reader may find interesting. Research has found significant dependencies between the tests Burciu and Simion (2019) and implementation issues with certain tests; for example, Hamano and Kaneko (2007) found that the settings of both the Discrete Fourier Transform test and Lempel-Ziv test were wrong, and Kowalska et al. (2022) found an error within the probability calculations for the Overlapping Template Matching test. Moreover, problems with how the results are analysed have been discovered; for example, Marton and Suciu (2015) found that although the NIST documentation provides guidance that the analysed RNG is random if all tests are passed, truly uniform data have a high probability (80%) of failing at least one NIST statistical (sub-)test. Some work has even suggested that the tests are “harmful” Saarinen , namely that “the weakest pseudo-random number generators will easily pass these tests, promoting false confidence in insecure systems”. During this work, we found an additional issue with the NIST test suite that we could not find reported elsewhere: the results showed that all tests failed whenever the CPU was being used for other computations simultaneously. The NIST Random Bit Generation Team have been made aware of this. Other test suites have also had their own reported problems, including Dieharder. In Sỳs et al. (2022), it was found that over 50% of the Dieharder tests generated biased null hypothesis distributions (which were expected to be uniform).
4 Statistical Testing of Different RNGs
@@
In this section, we use our $\mathsf{STE}$ to analyse the statistical properties of the random numbers produced by some commonly used RNGs. At this stage of our analysis, we do not apply any post-processing to the RNG’s output; however, some of the RNGs that we consider already have post-processing included, in the form of so-called conditioners or deterministic randomness extractors. Therefore, in cases where the post-processing already exists, our statistical analysis applies to the joint system, which comprises both the randomness (or entropy) source and the existing post-processing in the device. Additionally, we use and discuss a NIST min-entropy estimation tool, which provides a min-entropy estimate for our subsequent analysis incorporating various levels of post-processing.
The RNGs that we analyse are as follows:
  • 32-bit LFSR: a software pseudo-RNG;
@@
  • Intel RDSEED Jun and Kocher (1999): a hardware RNG based on thermal noise, i.e., a chaotic process;
@@
  • IDQuantique (IDQ) Quantis Quantique (2004): a hardware RNG based on the quantum effect of detecting photons at the output of a semi-transparent mirror.
@@
Further details and descriptions of the RNGs can be found in Section A.3. The LFSR remains widely used in numerical simulations, despite its well-known flaws Stępień and Walczak (2013). In this work, we primarily use it as a benchmark, serving as an example of a poor choice due to the patterns in its output and its short period. However, we will show that post-processing its output has a statistically significant impact. We note that IDQ’s Quantis is marketed as an RNG for cryptographic use, with certifications for compliance with various security and cryptographic standards. (IDQ Quantis has reportedly passed certifications or government validations, including “NIST SP800-22 Test Suite Compliance, METAS Certification, CTL Certification, multiple iTech Labs certificates, and compliance with BSI’s AIS31 standard (dedicated version of Quantis)” Quantique (2004)). These RNGs are used in numerous applications and are a sample of the different types of RNGs available today. The statistical analysis is performed using the run_all function in our statistical test environment on $10\times 10$ Gbit files from each RNG and, similarly, using the NIST test suite performed on $10\times 100$ Mbit files split into 100 sub-strings, each of $1$ Mbit. The NIST min-entropy estimators McKay et al. (2016) are used in the non-IID setting on $10\times 1$ Mbit files. This analysis far exceeds that required by certification bodies, so it may be a result of independent interest. All testing is performed using the default parameters, unless otherwise stated.
@@
An RNG producing near-perfect randomness should pass almost all statistical tests. More concretely, we mean that the ideal RNG would fail less than 7.5 of the 4600 individual statistical tests on average. (This number is the expected amount of type 2 error, i.e., the expected maximum number of failed tests, given that the underlying distribution is indistinguishable from uniform. Note that we are implicitly assuming that each statistical test is independent). The results obtained for the three RNGs are summarised in the following Table 3 and displayed visually in Figure 4 (level 0).
@@
Table 3: This table gives the average sum of statistical tests failed for $5\times 10$ Gbit samples from each RNG (after testing 10 samples). The results are presented in this way to allow for direct comparison to later results, where only $5\times 10$ Gbit samples are tested. Due to the 32-bit LFSR failing PractRand quickly, only 635 tests were conducted (instead of the full 4600), so we rescale these results. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses. The full results can be found in Appendix B.
{adjustwidth}
[…]
@@
In the statistical tests, the RDSEED RNG performs the best, failing the fewest tests, and the 32-bit LFSR performs the worst, failing the most. The poor performance of the LFSR is likely due to its periodicity, since bits repeat every $12^{32}-$ 1 ($4.3$ Gbits), and this is less than the size of the files tested. The IDQ Quantis device performs well in the NIST and Diehard tests but fails an ENT test and several tests in TestU01’s Rabbit and Alphabit suites. These observations reproduce (and add confidence to) the results of previous work Hurley-Smith and Hernandez-Castro (2020). These results, especially for IDQ’s device, exhibit the need to go beyond the requirements of certification bodies for statistical testing, with additional tests providing a noticeable advantage in detecting failures.
@@
The NIST min-entropy estimators McKay et al. (2016) are a collection of algorithms that give a standardised means of estimating the min-entropy (as defined in Figure 2) of an RNG’s output. These estimators are useful both in evaluating the entropy generation of the RNG and in calculating a min-entropy bound, which we later use to determine the parameters for randomness extractors. Although the estimators are designed to test entropy sources without post-processing, this is not feasible in the case of RDSEED or the IDQ Quantis device.
In Table 4, the NIST min-entropy estimator per byte is the average observed per-byte min-entropy calculated by the NIST min-entropy estimator tool, while $\overline{\mathsf{est}}$ is the per-bit average. The sample standard deviation, $\sigma$, reflects the variability in the different test results calculated using Equation 17, and $\alpha$ is a lower bound (with probability of at least $1-2^{-32}$) on the per-bit min entropy for any test sample. Details of this derivation can be found in Section B.2.1.
Table 4: This table shows the average NIST min-entropy estimator, the sample standard deviation, and a lower bound for min-entropy/bit for each RNG. These results are the average of 10 tests on different 1’000’000 bit samples, each generated with significant time gaps between the generation of each test sample. Full results can be found in Section B.2.
@@
The IDQ Quantis device had the highest estimated min-entropy per bit, with a value of 0.895, although all three RNGs had similar values. The 32-bit LFSR had the largest sample standard deviation, indicating the greatest fluctuation in the min-entropy estimates across different test samples. We note that the NIST SP800-90 series recommends min-entropy per bit of at least $1-2^{-32}$ for an RNG to be considered to have full entropy Buller et al. (2022). This value is significantly higher than any values that we observed. However, some of the NIST min-entropy estimator tests are known to produce significant underestimates Zhu et al. (2017), which potentially explains the large disparity between our estimates and the NIST requirement. Moreover, in our case, underestimates are not problematic, since we desire a lower bound on the min-entropy of the RNG’s output.
5 A Variety of Post-Processing Methods
Randomness extractors are mathematical algorithms that distil weakly random bit strings (more precisely, a necessary (but not sufficient) condition for randomness extraction to be successful is that the source has some min-entropy; see Figure 2), in the sense that they are not uniformly distributed, into a near-perfect random bit string. In this section, we present, implement, and test a variety of randomness extraction processes. The main question that we seek to answer is whether these methods have an observed impact on the statistical properties of the RNG’s output. The procedure that we follow is the following.